The value of x+y+Z =15 and 1/x+1/y+1/z is 5/3 .if a,x,y,z,b are in AP as well as are in HP then a and b are 

Grade:

1 Answers

Arun
5 years ago
Dear student
 
As we know
 
x + y + z = 3 (a + b)/2
 
Hence. a+ b = 10
 
Also a ,x ,y , z ,b are in Ap as well as HP . Hence all the numbwrs are same.
 
Hence
a = 5
 
Also b = 5